OpenStudy (anonymous):

Radicals: "Simplify. Assume all variable are positive." Problem below. I need help understanding *how* to solve it.

OpenStudy (anonymous):

\[\sqrt[3]{81x^2}\]

OpenStudy (anonymous):

I've gotten to \[\sqrt[3]{2^{3}\times9}\] but, afer that, I'm not sure. Is x^2 basically cancelled out or is it xtimesx?

OpenStudy (anonymous):

This is most likely wrong but if anyone could let me know.... the answer I got was \[2x \sqrt[3]{9x}\]

OpenStudy (cggurumanjunath):